The distance from Cleveland to Pittsburgh is about 2/5 of the distance from Cleveland to Chicago. About how far is Cleveland from Chicago? (It is 120 miles from Cleveland to Pittsburgh.) How do I set up an equation for this problem?
Answer by rfer(16322)

120/(2/5)=x
120*5/2=x
600/2=x
x=300